OpenStudy (anonymous):

Using Gaussian elimination, find all solutions to each of the given systems of equations expressing solutions as a linear combination of all vectors. 5x+2y+4z=24 2x-y+z=9 x-2y=z=8

OpenStudy (anonymous):

I worked out all of the Gaussian elimination and ended up with: x-2y+z=8 3y-z=-7 -z+16 I just don't know what to do from there

OpenStudy (unklerhaukus):

\[5x+2y+4z=24\\ 2x-y+z=9\\ x-2y\color{brown}\pm z=8\] \[\left[ \begin{array}{cc}5&2&4\\ 2&-1&1\\1&-2&\color{brown}\pm1\end{array}\right]\left[\begin{array}{c}x\\y\\z \end{array}\right] =\left[\begin{array}{c}24\\9\\8 \end{array}\right]\] \[\left[\begin{array}{c}5&2&4\\ 2&-1&1\\1&-2&\color{brown}\pm1 \end{array}\left|\begin{array}{c}24\\9\\8 \end{array}\right.\right]\]

OpenStudy (anonymous):

I already wrote the augmented matrix, as well as used Gaussian Elimination, I do not know how to write the solutions as a linear combination

OpenStudy (unklerhaukus):

dont stop eliminating until you get reduced row echelon form \[\left[\begin{array}{c}1&0&0\\ 0&1&0\\0&0&1 \end{array}\left|\begin{array}{c}x\\y\\z \end{array}\right.\right]\]

OpenStudy (anonymous):

That wasn't how I was taught. I am trying to make that equation that I have into a form like x -3 -2 y = -3 +t 5 z 0 1 as an example, I don't know how to draw matrices on here. I need to, through the gaussian elimation I used, turn my solutions into that form